When a ball is released freely, it falls to the ground at 25 degrees. What is the horizontal scale and maximum height of the ball?
1
Expert's answer
2021-01-19T07:09:50-0500
"\tan{25}=\frac{v_y}{v_x}=\frac{gt}{v_x}"

"\frac{x}{H}=\frac{v_xt}{0.5gt^2}=\frac{2}{\tan{25}}=4.3"